Suppose a certain traffic light shows a red light for 45 seconds, a green light for 45 seconds, and a yellow light for 5 seconds. If you look at this light at a randomly chosen instant, what is the probability that you see a red or yellow light? (Enter your probability as a fraction.)
Answer #1

Probability distribution of different colors of traffic light is as follows.

Colour Time Probability Red 45 9/19 Green 45 9/19 Yellow 5 1/19 Total 95 1

Required probability is given by

\tiny P\left ( \text {Red}\cup \text {Yellow} \right )=P\left ( \text {Red} \right )+P\left (\text {Yellow} \right )=\frac{9}{19}+\frac{1}{19}=\frac{10}{19}